# Enhancing LLM-based Test Generation for Hard-to-Cover Branches via Program Analysis

### v1

@@ III-B1 Backward Analysis
The goal of the backward analysis is to extract method-invocation sequences ending with the target method. These sequences represent diverse real usage scenarios of the target method in the project and have a high chance to capture the whole process of constructing objects used in the target method. For example, a method might create an object o and invoke the target method v with o as an argument. A backward method-invocation analysis would allow us to trace back to the object construction process.
@@
Call Graph Construction. A directed call graph is constructed to represent the method invocation relationships among methods in the module hosting the target method. The nodes in the graph represent the methods in the module extracted by traversing the abstract syntax tree (AST), and the edges represent invocation relationships between pairs of methods. Figure 4 illustrates an example of a constructed call graph. Given the simplified code snippet, the call graph will contain nodes method1 (denoted by ) from RelevantClass1, method2 (denoted by ) and method3 (denoted by ) from RelevantClass2. Clearly, is called by and , while is only called by .
For each extracted method, TELPA also records the associated classes (the parent class and classes utilized within the method), as well as their declarations and constructors. This information is important because extracted methods might belong to or interact with variables/methods from multiple different classes. Understanding where these elements come from is critical for LLMs to construct valid objects.
@@
Sequence Extraction. With the directed call graph, TELPA proceeds to extract sequences of method invocations for the target method. Specifically, TELPA identifies all the entry nodes from the graph and employs the Depth-First Search (DFS) strategy to gather all the paths originating from entry nodes and terminating at the target method. If a cycle is encountered during path extraction, the traversal of that path will stop to prevent creating infinite paths. It is worth noting that for all public methods inside a class, we consider each method itself as a complete path whose entry method is also the target method. Going back to the example in Figure 4, for the target method , three paths will be extracted: , and . As can be seen here, there might be multiples call paths from a same entry method to the target method (e.g., and ).
@@
To alleviate computational load, TELPA employs a straightforward filtering strategy to condense these sequences with the same entry method to the target method. That is, only the shortest path is kept. This strategy is adopted due to such a shorter and more direct call paths are easier for LLMs to understand and learn from. In our example, the path will be selected (that is, will be filtered out).
III-B2 Forward Analysis
@@
As illustrated in Section II, it is challenging for test generation techniques to reach certain branches when the outcome of the branch condition depends on other invoked methods. To address it, TELPA performs a forward method-invocation analysis to recursively collect all methods that might be relevant to determining which branch to execute. The extracted methods provide rich information for interpreting what really entails in branch conditions. Below, we describe the forward analysis process in detail with the example presented in Figure 5.
@@
To extract the set of all relevant methods for a target method (example in Figure 5), TELPA first identifies branch conditions and captures all variables (var1, param1) and methods (method2) these condition expressions contain. All involved methods (method2) are added to . For each involved variable, TELPA recursively extracts relevant assignment statements. In our example, the value of var1 is determined by method1, which takes an argument var0 determined by method0. TELPA considers both assignments are relevant to var1 and adds the involved methods (method0 and method1) to . In this way, all methods that can impact the outcome of branch conditions are retrieved and a precise search scope is identified to understand the behavior of condition expressions.
III-C Counter-Examples Sampling
@@
With the preceding testing process described in Section III-A, we have acquired a set of tests that fail to reach specific hard-to-cover branches. These existing tests can be used as counter-examples to guide LLMs to generate divergent tests, further enhancing the overall test efficiency. However, incorporating all existing generated tests into the prompt is impractical, which will not only impose significant computational burden on LLMs but also easily exceeds the input length limit of the prompt. To mitigate this, TELPA samples a diverse set of counter-example tests, striking a balance between effectiveness and efficiency. As our goal is to reach hard-to-cover branches, it is essential to let LLM understand which branches have already been covered by existing tests. Meanwhile, as mentioned before, we want to have a small (preferably minimal) set of tests as counter-examples.
@@
To achieve this, TELPA employs a coverage-based approach. For each method invocation sequence of a target method obtained from the backward analysis, TELPA collects all the existing tests that invoke the first method of the sequence (denoted by ) as the candidate tests (denoted by ). We only consider the first method of the sequence as it will in the end propagate to the target method through the chain of method calls. From , TELPA first selects the test that achieves the highest coverage for the target method , then picks the next tests that can achieve the highest incremental coverage. This is an iterative process until no more test can increase the coverage. In this way, the minimal set of tests can be acquired to reach all already-covered branches. As can be seen, the counter-examples are constructed for each method invocation sequence, this is because in the later stage of test generation, each time we will only feed one method invocation sequence to the LLM to alleviate the computational load and avoid exceeding the input length limit.
III-D Feedback-based Test Generation Process with LLMs
@@
TELPA incorporates program analysis results and counter-examples into the prompt for the LLM to generate new tests. Previous studies have shown large performance improvement for LLMs through Chain-of-Thought (CoT) [7, 8]. Therefore, we also adopt a typical CoT strategy and divide the process into two stages. Note that TELPA can also adopt different CoT techniques, as the contribution of TELPA lies in improving the prompting approach instead of designing CoT techniques.
@@
In the first stage, TELPA constructs the context of the target method with two sources: 1) methods in a selected method-invocation sequence from backward analysis, and 2) the associated methods from forward analysis, along with the declarations and constructors of their associated classes. To select the method-invocation sequence for in each iteration, TELPA first uses the sequence (i.e., the entry method is itself), as this is the most straightforward sequence for the LLM to invoke and understand the semantics of . If this sequence does not exist (e.g., the target method is not a public method), or the generated tests with it in the second stage do not cover new branches (note that our testing process is iterative with feedback), TELPA randomly selects an unused sequence. The context information is then fed to the LLM for summarizing the functionality of , enabling the LLM to gain deep understandings of the method semantics. The test generation process for terminates when no sequence is left.
@@
In the second stage, TELPA integrates the contents of the sampled counter-examples corresponding to the method invocation sequence into the prompt, along with a description: “these counter-examples enter the target method via the selected sequence of method invocations.” Such a description can facilitate the understanding of the LLM on the intent of the generation, namely to cover different branches in the target method following the method invocation sequence. The LLM is then instructed to generate divergent tests.
@@
The new test generated by the LLM is executed and added to the existing tests along with its coverage information. The used counter-examples are removed to ensure the diversity in future counter-example selection. This type of feedback helps TELPA generate more tests with different behaviors.
IV Evaluation
@@
RQ1: To what extent can TELPA improve code coverage compared to the state-of-the-art techniques? The goal of TELPA is to reach hard-to-cover branches, and this RQ verifies its ability in this regard.
RQ2: How does each of the main components in TELPA affect the performance? This RQ leads to an ablation study to examine how TELPA’s main components (i.e., method-invocation analysis, counter-example sampling, coverage-based feedback) impact the performance.
@@
RQ3: How does the adoption of preceding test generation tools and LLMs impact the performance of TELPA? Our TELPA outsources initial test generation to existing tools, and utilizes an LLM to generate new tests. In this RQ, we examine how the adoption of existing tools and the choice of the LLM can impact the performance.
IV-A Experimental Setup
IV-A1 Implementation of TELPA
While TELPA is not limited to specific programming languages, in this experiment, we target test generation for Python programs, given the popularity of Python and the existing Python benchmarks [4]. TELPA is also developed in Python 3.10 with over 6k lines of code. The LLM Phind-CodeLlama-34B-v2 [9] from Hugging Face [10] is employed for test generation, as it is one of the most effective open-source code LLMs. PyTorch 1.13 [11], Hugging Face Transformers 4.35.2 [12], and fastchat 0.2.30 [13] are used to run the LLM locally. Following the existing work [14, 15], we set the temperature to 1, enabling TELPA to obtain results from the LLM with a high degree of confidence.
@@ IV-A1 Implementation of TELPA
To reduce the resource consumption, TELPA outsources initial test generation to state-of-the-art techniques. In our study, we adopted two test generation tools separately: a state-of-the-art SBST technique Pynguin [1] and a state-of-the-art LLM-based technique CODAMOSA [4]. That is, we constructed two instances of TELPA: Pynguin-based TELPA and CODAMOSA-based TELPA. The adoption of these two different types of preceding test generation techniques ensures the conclusion generalizability. More information regarding Pynguin and CODAMOSA will be given in Section IV-A2.
@@
As mentioned before, TELPA is activated when the preceding tool fails to increase coverage within a pre-defined timeframe. As the previous work [4], we conducted a pilot study on a small test benchmark to optimize the performance of TELPA. The timeframe is set to 2min based on the observed results, namely TELPA will take over the test generation task if Pynguin or CODAMOSA cannot increase the coverage for two minutes. This choice is further discussed in Section IV-E.
IV-A2 Baselines for RQ1
@@
To answer RQ1, we compared TELPA with Pynguin and CODAMOSA, the techniques used for initial test generation in TELPA.
@@
Pynguin [1] is the state-of-the-art SBST technique for dynamically typed programming languages like Python. It initiates the testing process by taking the Python code under test as input, and employs one of search-based algorithms (including MIO [16, 17], MOSA [18], and DynaMOSA [18]) to generate new tests by mutating values and statements of a seed test. Here, we used Pynguin 0.19.0 with DynaMOSA as this algorithm has been demonstrated to be the most effective one among all supported algorithms in Pynguin [1].
@@
CODAMOSA [4] is the state-of-the-art LLM-based test generation tool, which was proposed for improving SBST techniques. It prompts LLMs to generate seed tests for Pynguin in order to help SBST escape the local optimum. More specifically, when Pynguin falls into the local optimum, CODAMOSA switches to invoke LLMs to generate new seed tests by incorporating as much source code of the module under test into the prompt as possible, and then switches back to Pynguin for running on these new seed tests, until the given testing budget is reached. CODAMOSA originally employed Codex as the LLM for generating seed tests. However, as the Codex service has been shut down [19], we employed Phind-CodeLlama-34B-v2 as the LLM for CODAMOSA to have a fair comparison with our technique.
One of the main contributions of TELPA is the novel prompting method. Unlike CODAMOSA that incorporates as much source code of the module under test as possible for prompting, TELPA only feed part of the code which is highly relevant to branch constraints to LLMs. The comparison with the LLM-based CODAMOSA cannot clearly demonstrate the impact of this new prompting method due to the different LLM invocation workflows. Therefore, we constructed a new LLM-based baseline (called BasicPrompt), which replaces our novel prompting method with the one used CODAMOSA (i.e., incorporating as much code as possible). In fact, this traditional prompting approach is commonly used to generate tests in practice [4, 5]. Since in our study, both Pynguin and CODAMOSA are used as preceding test generation tools, we created two instances of BasicPrompt for comparisons: Pynguin-based BasicPrompt and CODAMOSA-based BasicPrompt.
@@
To sum up, we have two sets of comparisons: 1) TELPA vs. Pynguin vs. BasicPrompt and 2) TELPA vs. CODAMOSA vs. BasicPrompt. These comparisons aim to investigate whether TELPA can improve the existing test generation tools (Pynguin and CODAMOSA) and examine whether our novel prompting method can outperform the widely-used prompting method regardless of the preceding test generation tools.
In our study, we did not compare TELPA with the other LLM-based test generation techniques such as ChatUnitTest [20] and ChatTester [5]. The reason is that these approaches leveraged similar prompting method as CODAMOSA. That is, they also feed LLMs with as much available information as possible, such as target methods and other methods around target methods. Particularly, their main novelty which contributes the performance improvement is the post-processing strategies adopted to repair invalid tests generated by LLMs. Those test-repair strategies are orthogonal to the contribution of TELPA, which can be also integrated into TELPA to further improve its overall test effectiveness.
IV-A3 Variants for RQ2

### v2

@@ 3.2.1. Object Construction Analysis
The goal of the object construction analysis is to extract method-invocation sequences ending with the target method. These sequences represent diverse real usage scenarios of the target method in the project and have a high chance to capture the whole process of constructing objects used in the target method. For example, a method may create an object o and invoke the target method v with o as an argument. A object construction analysis could collect the invocation sequence v and trace back to the object construction process.
@@
Call Graph Construction. TELPA constructs a directed call graph to model method invocation relationships within the module containing the target method. The nodes in the graph represent the methods in the module, and the edges represent invocation relationships between pairs of methods. Figure 5 illustrates a simplified version of the code shown in Figure 4 and the constructed call graph. Given the simplified code snippet, the call graph will contain nodes method1 (denoted by ) and method2 (denoted by ) from Class1, and method3 (denoted by ) from Class2. Clearly, is called by and , while is only called by .
@@
Sequence Extraction. With the directed call graph, TELPA proceeds to extract sequences of method invocations for the target method. Specifically, Depth-First Search (DFS) is used to identify all paths starting from entry nodes and terminating at the target method. If a cycle is encountered during path extraction, the traversal of that path will stop to prevent creating infinite paths. Going back to the example in Figure 5, for the target method , three paths will be extracted: , and .
@@
To alleviate computational load, only the shortest path from each entry point to the target method is retained. This strategy is used as such shorter and more direct call paths are easier for LLMs to understand and learn from. In our example, the path is selected (that is, is filtered out).
3.2.2. Branch Dependency Analysis
@@
As illustrated in Section 2, it is challenging for test generation techniques to reach certain branches when the outcome of the branch condition depends on other invoked methods. To address it, TELPA performs a branch dependency analysis to recursively collect all methods that might be relevant to determining which branch to execute. The extracted methods provide rich information for interpreting what really entails in branch conditions. Below, we describe the process in detail with the example presented in Figure 6.
@@
To extract the set of associated methods for a target method (method1 in Figure 6), TELPA first identifies branch conditions and captures all variables (var0) and methods (dependent_method0) that these condition expressions contain. All involved methods (dependent_method0) are added to . For each involved variable, TELPA recursively extracts relevant assignment statements. In our example, the value of var0 is determined by dependent_method1. Therefore, TELPA considers dependent_method1 is relevant to var0 and adds the involved methods (dependent_method1) to . In this way, all methods that can impact the outcome of branch conditions are retrieved and a precise search scope is identified to understand the behavior of condition expressions.
Note that TELPA does not always use both branch dependency analysis and object construction analysis simultaneously. Specifically, it applies these analyses selectively based on the characteristics of the target method. Branch dependency analysis is used when the target method involves intricate inter-procedural dependencies in branch conditions, and object construction analyses is used when the target method involves complex objects in branch conditions. If the target method involves both complex objects and inter-procedural dependencies, TELPA applies both analyses to address these challenges comprehensively. For the simpler cases where the target method does not involve complex objects and inter-procedural dependencies, TELPA provides the method under test itself as its context to the LLM, like the existing LLM-based approaches (such as CHATTESTER). This ensures that TELPA remains effective for such cases.
3.3. Counter-Examples Sampling
@@
With the preceding testing process described in Section 3.1, we have acquired a set of tests that fail to reach specific hard-to-cover branches. These existing tests can be used as counter-examples to guide LLMs to generate divergent tests, further enhancing the overall test efficiency. However, incorporating all existing generated tests into the prompt is impractical, which will not only impose significant computational burden on LLMs but also easily exceed the input length limit of the prompt. To mitigate this, TELPA samples a diverse set of counter-example tests, striking a balance between effectiveness and efficiency. As our goal is to reach hard-to-cover branches, it is essential to let LLM understand which branches have already been covered by existing tests. Meanwhile, as mentioned before, we want to have a small (preferably minimal) set of tests as counter-examples.
@@
To achieve this, TELPA employs a coverage-based approach. For each method invocation sequence of a target method obtained from the object construction analysis, TELPA collects all the existing tests that invoke the first method of the sequence (denoted by ) as the candidate tests (denoted by ). We only consider the first method of the sequence as it will in the end propagate to the target method through the chain of method calls. From , TELPA first selects the test that achieves the highest coverage for the target method , then picks the next tests that can achieve the highest incremental coverage. This is an iterative process until no more test can increase the coverage. In this way, the minimal set of tests can be acquired to reach all already-covered branches. Note that the counter-examples are constructed for each method invocation sequence, this is because in the later stage of test generation, each time we will only feed one method invocation sequence to the LLM to alleviate the computational load and avoid exceeding the input length limit.
3.4. Feedback-based Test Generation Process with LLMs
@@
TELPA integrates program analysis results and counter-examples into prompts for the LLM to generate new tests. Previous studies have shown large performance improvement for LLMs through Chain-of-Thought (CoT) (Wei et al., 2022; Zhang et al., 2022). Therefore, we also adopt a typical CoT strategy and divide the process into two stages. Note that TELPA can also adopt different CoT techniques, as the contribution of TELPA lies in improving the prompting approach instead of designing CoT techniques.
@@
In the first stage, TELPA constructs the context of the target method with two sources: 1) methods in a selected method-invocation sequence from object construction analysis, and 2) the associated methods from branch dependency analysis. For each method, TELPA incorporates the declarations and constructors of their associated classes and the content of the method itself as the context information. To select the method-invocation sequence for in each iteration, TELPA first uses the sequence (i.e., the entry method is itself), as this is the most straightforward sequence for the LLM to invoke and understand the semantics of . If this sequence does not exist (e.g., the target method is not a public method), or the generated tests with it in the second stage do not cover new branches (note that our testing process is iterative with feedback), TELPA randomly selects an unused sequence. The context information is then fed to the LLM for summarizing the functionality of , enabling the LLM to gain contextual awareness of method semantics. The test generation process for terminates when no sequence is left. For example, given method1 in Figure 4, if the selected sequence is , the prompt will be constructed using method3 from the object construction analysis, dependent_method0 and dependent_method1 from the branch dependency analysis, and method1 itself, as shown in Figure 7.
@@
In the second stage, TELPA integrates the contents of the sampled counter-examples corresponding to the method invocation sequence into the prompt, along with a description: “These counter-examples enter the target method via the selected sequence of method invocations. They can cover different parts of the method. Please generate new test cases that cover different scenarios or edge cases. ”, as shown in Figure 7. Such a description can facilitate the understanding of the LLM on the intent of the generation, namely to cover different branches in the target method following the method invocation sequence. The LLM is then instructed to generate divergent tests.
@@
The new test generated by the LLM is executed and added to the existing tests along with its coverage information. The used counter-examples are removed from the candidates to ensure the diversity in future counter-example selection. This type of feedback helps TELPA generate more tests with different behaviors.
4. Evaluation
@@
RQ1: To what extent can TELPA improve code coverage compared to the state-of-the-art techniques?
RQ2: How does each of the main components in TELPA affect the performance?
@@
RQ3: How does the adoption of preceding test generation tools and LLMs impact the performance of TELPA?
RQ4: How is the correctness of the unit tests generated by TELPA?
RQ5: How generalizable is TELPA?
4.1. Experimental Setup
4.1.1. Implementation of TELPA
While TELPA is not limited to specific programming languages, in this experiment, we target test generation for Python programs in RQ1-RQ4, given the popularity of Python and the existing Python benchmarks (Lemieux et al., 2023). Additionally, we adapted TELPA to Java to evaluate its generalizability in RQ5.
Regarding the Python version of TELPA, to construct the call graph, we rely on PyCG (Salis et al., 2021), a state-of-the-art call graph generator for Python. PyCG provides a robust foundation for identifying method invocation relationships within the code. PyCG constructs an assignment graph that tracks the assignment relations between program identifiers, such as variables and classes. Furthermore, it uses namespace and scope resolution to distinguish between methods with the same name in different classes. To better align with TELPA’s requirements for sequence extraction and prompt construction, we map the nodes in the PyCG-extracted call graph back to the corresponding AST nodes parsed using Python’s AST module. Since TELPA focuses on generating tests for the target module, we primarily consider method invocations within the module under test. External functions (e.g., those from imported libraries) are not further analyzed for sequence extraction, as their internal implementations are outside the scope of the target module. For coverage analysis, TELPA uses coverage.py (cov, 2024), a widely-used tool for measuring code coverage in Python programs. The LLM Phind-CodeLlama-34B-v2 (phi, 2024) from Hugging Face (hug, 2024a) is employed for test generation by default, as it is one of the most effective open-source code LLMs according to the LLM leaderboard maintained by Hugging Face (lea, 2024). PyTorch 1.13 (tor, 2024), Hugging Face Transformers 4.35.2 (tra, 2024), and fastchat 0.2.30 (fas, 2024) are used to run the LLM locally. Following the existing work (Xia et al., 2023; Yang et al., 2023), we set the temperature to 0, enabling TELPA to obtain more deterministic results from the LLM.
@@ 4.1.1. Implementation of TELPA
To reduce the resource consumption, TELPA uses state-of-the-art techniques for initial test generation. In our study, we adopted two test generation tools separately: a state-of-the-art SBST tool Pynguin (Lukasczyk and Fraser, 2022) and a state-of-the-art LLM-enhanced technique CODAMOSA (Lemieux et al., 2023). That is, we constructed two instances of TELPA: Pynguin-based TELPAp and CODAMOSA-based TELPAc. The adoption of these two different types of preceding test generation techniques ensures the conclusion generalizability. More information regarding Pynguin and CODAMOSA will be given in Section 4.1.2.
Specifically, to answer RQ5, we adapted TELPA to Java. Note that although TELPA is in theory language-independent, it still requires engineering efforts as the analysis and the parsing of test cases are highly specific to the chosen frameworks. Thus we made the engineering effort to extend TELPA to support Java projects. Specifically, we adapted TELPA’s program analysis components to work with Java code by incorporating JavaParser (jav, 2024), allowing to effectively parse Java code. We also implemented the test execution process using the JUnit framework (the standard framework for unit testing in Java) and incorporated Jacoco (jac, 2024) for code coverage analysis (a widely-used tool for measuring Java code coverage). Additionally, we modified the processing of the LLM’s output to extract Java tests and address potential issues such as missing external libraries.
Similarly, to reduce resource consumption, TELPA leverages the state-of-the-art search-based test generation tool for Java (i.e., EvoSuite (Fraser and Arcuri, 2011b)) for preceding test generation. Specifically, EvoSuite is used to generate initial tests, when EvoSuite no longer improves coverage within a set time window, TELPA is activated, takes over the test generation process entirely and leverages EvoSuite-generated tests as counter-examples. The LLM is used in the same way as the Python version of TELPA.
@@
As mentioned before, TELPA is activated when the preceding tool fails to increase coverage within a pre-defined timeframe. As previous work (Lemieux et al., 2023), we conducted a pilot study on a small test benchmark to optimize the performance of TELPA. The timeframe is set to 2 minutes based on the observed results, namely TELPA will take over the test generation task if Pynguin, CODAMOSA or EvoSuite cannot increase the coverage for two minutes. This choice is further discussed in Section 4.7.
4.1.2. Baselines for RQ1 and RQ4
@@
To answer RQ1 and RQ4, we compared TELPA with three baselines: Pynguin, CODAMOSA, and CHATTESTER.
@@
Pynguin (Lukasczyk and Fraser, 2022) is the state-of-the-art SBST tool for Python. It initiates the testing process by taking the Python code under test as input, and employs search-based algorithms (including MIO (Arcuri, 2017, 2018), MOSA (Panichella et al., 2017), and DynaMOSA (Panichella et al., 2017)) to generate new tests by mutating values and statements of a seed test. Here, we used Pynguin 0.19.0 with DynaMOSA as this algorithm has been demonstrated to be the most effective one among all supported algorithms in Pynguin (Lukasczyk and Fraser, 2022).
@@
CODAMOSA (Lemieux et al., 2023) is the state-of-the-art LLM-based test generation tool, which was proposed for improving SBST techniques. It prompts LLMs to generate seed tests for Pynguin in order to help SBST escape the local optimum. More specifically, when Pynguin falls into the local optimum, CODAMOSA switches to invoke LLMs to generate new seed tests by incorporating as much source code of the module under test into the prompt as possible, and then switches back to Pynguin for running on these new seed tests, until the given testing budget is reached. Specifically, we consider CODAMOSA as an integral tool integrating both Pynguin and the LLM, and then incorporate it into the preceding test generation component of TELPA. In the case of using CODAMOSA as the preceding test generation tool, when the switch (between Pynguin and the LLM) within CODAMOSA does not improve the coverage any more within the given timeframe (exceeding the switching time used within CODAMOSA), TELPA makes CODAMOSA terminate and then activates its program-analysis-enhanced LLM-based test generation component for further coverage improvement. CODAMOSA originally employed Codex as the LLM for generating seed tests. However, as the Codex service has been shut down (cod, 2024), we employed Phind-CodeLlama-34B-v2 as the LLM for CODAMOSA to have a fair comparison with our technique. We call it CODAMOSA.
CHATTESTER (Yuan et al., 2024) is another state-of-the-art LLM-based test generation tool, which first extracts context information for ChatGPT to generate tests and then designs post-processing strategies to fix invalid tests. The main contribution of TELPA is the novel prompting method. Unlike CODAMOSA that incorporates as much source code of the module under test as possible for prompting, or CHATTESTER that incorporates insufficient context for prompting, TELPA feeds the code information highly relevant to branch constraints to LLMs. However, CODAMOSA and TELPA differ in both their LLM invocation workflows and their prompting methods. Specifically, CODAMOSA iteratively switches between the search-based tool Pynguin and the LLM, while once TELPA switches, it does not revert back to the preceding tool. That is, TELPA continues to use the LLM with its program-analysis-enhanced prompting for subsequent test generation. Additionally, CODAMOSA uses a simple prompting method by incorporating the entire module’s source code as context for the LLM. TELPA introduces a novel prompting method that leverages program analysis to provide precise and relevant context for the LLM. Therefore, the comparison with the LLM-enhanced CODAMOSA cannot clearly demonstrate the impact of this new prompting method due to the different LLM invocation workflows. Therefore, we used CHATTESTER as another LLM-based baseline. To highlight the effectiveness of advanced prompting methods for test generation, we directly replaced our novel prompting method with the one used by CHATTESTER and removed the post-processing strategy from CHATTESTER. Note that this component is orthogonal to the test generation process and can be also applied to TELPA. That is, we activated the prompting method of CHATTESTER for test generation when the preceding test generation tool reaches a coverage plateau, same as the workflow of TELPA. CHATTESTER originally employed ChatGPT as the LLM for generating tests, we replaced it with Phind-CodeLlama-34B-v2 for fair comparison. We call it CHATTESTER. Since both Pynguin and CODAMOSA are used as preceding test generation tools in our study, we also created two instances of CHATTESTER for comparisons: Pynguin-based CHATTESTER and CODAMOSA-based CHATTESTER.
@@
To sum up, we have two sets of comparisons: 1) TELPAp vs. Pynguin vs. CHATTESTER and 2) TELPAc vs. CODAMOSA vs. CHATTESTER. These comparisons aim to investigate whether TELPA can improve the existing test generation tools (Pynguin and CODAMOSA) and examine whether our novel prompting method can outperform the state-of-the-art prompting method proposed by CHATTESTER regardless of the preceding test generation tools.
4.1.3. Variants for RQ2